SOlve the equation
3(n-1)^2+20=19(n-1)

Answer by fractalier(6550)

Okay from
3(n-1)^2 + 20 = 19(n-1)
we expand and get
3n^2 - 6n + 3 + 20 = 19n - 19
now collect terms and solve, we get
3n^2 - 25n + 42 = 0
(3n - 7)(n - 6) = 0
n = 7/3 or n = 6